# Masked Attribute Description Embedding for Cloth-Changing Person Re-identification

### v2

@@ 4.1 Datasets
4.2 Attribute Analysis
@@
As depicted in Fig. 3, the figure illustrates the higher maintained proportions of cloth-irrelevant attributes in each dataset. It is observed that the biological attributes (age, gender, hair) and carrying objects of persons generally do not change across cameras in the short term. Additionally, features like shoe type and color generally stabilize even as persons change their clothing. These attributes can be important in the model learning clothing-insensitive features.
Refer to caption (a) PRCC Refer to caption (b) LTCC Refer to caption (c) Celeb-reID-light Refer to caption (d) LaST
Figure 3: Retain ratio of irrelevant attributes of persons’ clothing in each dataset.
4.3 Evaluation Settings and Protocol
For the cloth-changing datasets PRCC [44] and LTCC [31], we follow their respective evaluation protocols and evaluate the performance under the clothes-changing and standard settings.
4.4 Implementation Details
@@
The input images are resized to 224×224 for all datasets. We divide the 24 layers of EVA02-large  [12] into three stages, and the number of layers of the Trv blocks is 10, 10, and 4, respectively(We discussed in 4.6.2). We use LayerNorm  [1] to normalize features. For data augmentation, we employ random cropping and random erasing [51]. Due to the limit of GPU memory, the batchsize is set to 8, each batch includes two different people, and the number of images for each person is 4. The SGD optimizer is employed in the optimization process, and 60 epochs are required. Moreover, the weight decay for the experiment is $5e^{-2}$. The warmup learning rate is initially set to $7.8125e^{-7}$. The learning rate is initially set to $2e^{-5}$ and divided by 100 at 40 and 60 epochs. The optimal parameter values are directly used for the other datasets without tuning.
4.5 Comparison with State-of-the-Art
@@
For RRCC and LTCC, we combine our proposed MADE method with some cloth-changing re-identification methods (i.e., SPT+ASE [44], GI-ReID [23], CESD [31], RCSANet [21], 3DSL [4], FSAM [19], BSGA+CRE [30], CAL [15], DCR-ReID [8], CCFA [16], AIM [45] and chan et al. [3]) were compared. We compare Celeb-reID-light with four cloth-changing re-identification methods (RCSANet [21], MBUNet [47], IRANet [36], and DeSKPro [41]) and some traditional methods. We compare LaST with CAL and some traditional methods. It is worth noting that among these CC-ReID methods, SPT+ASE, GI-ReID, CESD, 3DSL, FSAM, BSGA+CRE, and DCR-ReID all integrate person multi-modal biometric features into the model to remove clothing interference. CAL and AIM mine the information of original RGB images. CCFA adopts the feature enhancement method. The method proposed by Chen et al. is based on the GAN network.
@@
Table 1: Evaluations on the PRCC and LTCC datasets (%), where ”sketch,” ”pose,” ”sil.”, ”parsing” and ”3D” denote contour sketches, keypoints, silhouettes, human parsing, and 3D shape information. Bold and underlined numbers are the top two scores.
@@
Method Venue Modality PRCC LTCC CC SC CC Genral rank-1 mAP rank-1 mAP rank-1 mAP rank-1 mAP SPT+ASE  [44] TPAMI 19 Sketch 34.4 - 64.2 - - - - - CESD  [31] ACCV 20 RGB+pose - - - - 26.1 12.4 71.4 34.3 RCSANet  [21] ICCV 21 RGB 48.6 50.2 100.0 97.2 - - - - 3DSL  [4] CVPR 21 RGB+pose+sil.+3D - 51.3 - - 31.2 14.8 - - FSAM  [19] CVPR 21 RGB+pose+sil. 54.5 - 98.8 - 38.5 16.2 73.2 35.4 GI-ReID  [23] CVPR 22 RGB+sil. - 37.5 - - 23.7 10.4 63.2 29.4 BSGA+CRE  [30] BMVC 22 RGB+parsing 61.8 58.7 99.6 97.3 - - - - CAL  [15] CVPR 22 RGB 55.2 55.8 100.0 99.8 40.1 18.0 74.2 40.8 CCFA  [16] CVPR 23 RGB 61.2 58.4 99.6 98.7 45.3 22.1 75.8 42.5 AIM  [45] CVPR 23 RGB 57.9 58.3 100.0 99.9 40.6 19.1 76.3 41.1 DCR-ReID  [8] TCSVT 23 RGB+parsing 57.2 57.4 100.0 99.7 41.1 20.4 76.1 42.3 chan et al. [3] ACM 23 RGB 58.4 58.6 100.0 99.7 32.9 15.4 73.4 36.9 MADE RGB+description 67.5 64.2 100.0 98.9 46.9 25.0 82.2 49.3
@@
Results on PRCC. We compare our method with twelve cloth-changing re-identification methods on PRCC in Table 1. We can notice that our method outperforms all other methods in the clothes-changing setting. Compared with AIM, the method of mining original RGB images in the clothes-changing setting, the rank-1 of our method increased by 9.6%, and the mAP increased by 5.9%. These shows that using multi-modal attribute description in CC-ReID to assist re-identification effectively improves results. Compared with the best method using multi-modal biometrics, BSGA+CRE, in the clothes-changing setting, the rank-1 increased by 5.7%, and the mAP increased by 5.5%. It shows that our method uses editable multi-modal attribute information and has better results when it is more convenient to remove clothing information than biological information. Data augmentation can significantly improve the model improvement effect, and our method is even better than CCFA, which uses feature enhancement. In the clothes-changing setting, the rank-1 increased by 6.3%, and the mAP increased by 5.8%.
@@
Results on LTCC. We compare our method with eight cloth-changing re-identification methods on LTCC in Table 1. Our method outperforms all other methods in both clothes-changing and general settings. Compared with CCFA, in the clothes-changing setting, The rank-1 increased by 1.6%, and the mAP increased by 2.9%. In the general setting, The rank-1 increased by 6.4%, and the mAP increased by 6.8%. Compared with AIM, in the clothes-changing setting, The rank-1 increased by 6.3%, and the mAP increased by 5.9%. In the general setting, The rank-1 increased by 5.9%, and the mAP increased by 8.2%. Compared with chan et al., in the clothes-changing setting, The rank-1 increased by 14.0%, and the mAP increased by 9.6%. In the general setting, The rank-1 increased by 8.8%, and the mAP increased by 12.4%. It shows that our method embedding fused image features with attribute description is better than using GAN networks to mine clothing-independent features of original images.
@@
Table 2: Evaluations on Celeb-reID-light(%). Bold and underlined numbers are the top two scores.
@@
Method Type Method Venue Celeb-reID-light rank-1 mAP Traditional OSNet [52] ICCV 19 21.3 11.7 DG-Net [50] CVPR 19 23.5 12.6 BoT(resnet50) [29] CVPRW 19 24.2 13.6 AGW(resnet50_nl) [46] TPAMI 21 30.2 15.4 TransReID [17] ICCV 21 31.3 18.6 CC-ReID RCSANet  [21] CVPR 21 29.5 16.7 MBUNet  [47] ICME 22 33.9 21.3 IRANet  [36] IVC 22 46.2 25.4 DeSKPro  [41] ICIP 22 52.0 29.8 MADE 69.6 52.1
Table 3: Evaluations on LaST(%). Bold and underlined numbers are the top two scores.
@@
Method Type Method Venue Celeb-reID-light rank-1 mAP Traditional OSNet [52] ICCV 19 64.3 21.0 BoT [29] CVPRW 19 67.1 23.6 HOReID [40] CVPR 20 68.3 25.5 Top-DB-Net [32] ICPR 20 69.4 25.0 CtF [39] ECCV 20 70.0 26.5 CC-ReID CAL [15] CVPR 22 73.7 28.8 MADE 82.6 45.8
@@
Results on Celeb-reID-light and LaST. We compare our method with some traditional and cloth-changing re-identification methods on the two datasets in Table 2 and Table 3. Our method outperforms all previous methods. The face is the most direct information for re-identification. For Celeb-reID-light, our method is even better than DeSKPro, which uses facial features. The rank-1 increased by 17.6%, and the mAP increased by 22.3%. For LaST, our method is better than CAL. The rank-1 increased by 8.9%, and the mAP increased by 17.0%. LaST is a large and challenging dataset that requires high model training complexity. Currently, there are few CC-ReID methods tested using LaST. It shows that our method is effective, has low model complexity, and can achieve testing of large datasets.
4.6 Ablation Study
@@
We use EVA02-large  [12] encoded by cloth label embedding  [34] as the baseline and also use the loss function in 3.4 for supervision. In this section, we explore the role of masked attribute description on the model’s learning of clothes-irrelevant features and the impact of different layer numbers in EVA02-large.
4.6.1 Effectiveness of Description
@@
We embed the complete attribute description vector (attribute description) and the attribute description vector after mask clothing information (masked attribute description) into the baseline, respectively, and summarize the experimental results in table4. To verify the effectiveness of attribute description in improving re-identification results. We conducted experiments on the clothes-changing setting of PRCC and LTCC. In both datasets, the performance of adding attribute description and masked attribute description is almost higher than the baseline.
The results of adding attribute description are higher than the baseline, indicating that even after persons change clothes, clothing-related attributes still play an essential role in the model. After masking the clothing information in the attribute description and embedding, for PRCC, the Rank-1 increased by 5.0%, and the mAP increased by 5.9%. For LTCC, the Rank-1 increased by 0.7%, and the mAP increased by 1.1%. The model is more accurate in person re-identification, indicating that this method can compel the model to discard clothing information and learn clothing-insensitive features.
@@
Table 4: Ablation studies of attribute description of MADE in clothes-changing setting on PRCC and LTCC.
Method PRCC LTCC rank-1 mAP rank-1 mAP baseline 62.5 58.3 46.2 23.9 baseline + attribute 66.3 62.7 45.4 24.2 baseline + masked attribute(MADE) 67.5 64.2 46.9 25.0
4.6.2 Number of Stages and TrV Blocks

### v3

@@ IV-A Datasets
For the cloth-changing datasets PRCC [1] and LTCC [17], we follow their respective evaluation protocols and evaluate the performance under the cloth-changing and standard settings. In Figure 4, we present examples of the datasets this paper used.
Figure 4: Examples of the datasets this paper used.
IV-B Person Attribute Analysis
@@
According to Fig. 5, we can observe that for PRCC and LTCC, which are both collected from real-world scenarios with short data collection periods and fixed scopes, the biological attributes of pedestrians are maintained at a relatively high proportion. Due to the challenging style of the LTCC for attribute recognition models, in the ablation experiments of LTCC (section IV-E1), the addition of the DEM module only marginally improves recognition accuracy. We discuss the impact of attribute recognition model accuracy on the experimental results in section IV-E1. Surprisingly, for Celeb-ReID-light and LaST, which originate from internet images, their biological attribute features are also maintained at a high proportion. In Fig. 5, we compute the average retention ratio for the four datasets. It can be observed that biological attributes such as age, gender, and hairstyle usually remain unchanged in the short term across different cameras. Additionally, even if individuals change clothes, features such as shoe type and color typically remain stable. These findings suggest that these attributes may be crucial for models to learn cloth-agnostic features, indicating that leveraging these stable attributes could aid models in better understanding and identifying individuals regardless of clothes variations.
Figure 5: Retain ratio of clothes irrelevant person attributes in each dataset. (a) PRCC, (b) LTCC, (c) Celeb-reID-light, (d) LaST, and (e) Average statistic.
IV-C Implementation Details
@@
The input images are resized to 224$\times$224 for all datasets. We divide the 24 layers of EVA02-large [13] into three stages, and the number of layers of the Trv blocks is 10, 10, and 4, respectively(We discussed in section IV-E3). We use LayerNorm [37] to normalize features. For data augmentation, we employ random cropping and random erasing  [38]. Due to the limit of GPU memory, the batch size is set to 8, each batch includes two different people, and the number of images for each person is 4. The SGD optimizer is employed in the optimization process, and 60 epochs are required. Moreover, the weight decay for the experiment is $5e^{-2}$. The warmup learning rate is initially set to $7.8125e^{-7}$. The learning rate is initially set to $2e^{-5}$ and divided by 100 at 40 and 60 epochs. The optimal parameter values are directly used for the other datasets without tuning.
IV-D Experimental Comparison
@@
For RRCC and LTCC, we combine our proposed MADE method with some cloth-changing re-identification methods (i.e., SPT+ASE [1], GI-ReID [5], CESD [17], RCSANet [39], 3DSL [3], FSAM [2], BSGA+CRE [15], CAL [9], DCR-ReID [40], CCFA [41], AIM [42] and chan et al. [10]) were compared. We compare Celeb-reID-light with four cloth-changing re-identification methods (RCSANet [39], MBUNet [4], IRANet [43], and DeSKPro [32]) and some traditional methods. We compare LaST with CAL and some traditional methods. It is worth noting that among these CC-ReID methods, SPT+ASE, GI-ReID, CESD, 3DSL, FSAM, BSGA+CRE, and DCR-ReID all integrate person multi-modal biometric features into the model to remove clothes interference. CAL and AIM mine the information of original RGB images. CCFA adopts the feature enhancement method. The method proposed by Chen et al.. is based on the GAN network. Considering the accuracy of the attribute extraction model, our experimental results are obtained under the premise of introducing 10% random 0-1 noise into the masked-attribute description. Discussions on the accuracy of the attribute extraction model and the proportion of noise can be found in Section IV-E1.
@@
TABLE I: Evaluations on the PRCC and LTCC datasets (%), where ”sketch,” ”pose,” ”sil.”, ”parsing” and ”3D” denote contour sketches, keypoints, silhouettes, human parsing, and 3D shape information. Bold and underlined numbers are the top two scores.
@@
Method Venue Modality PRCC LTCC CC SC CC Genral rank-1 mAP rank-1 mAP rank-1 mAP rank-1 mAP SPT+ASE  [1] TPAMI 19 Sketch 34.4 - 64.2 - - - - - CESD  [17] ACCV 20 RGB+pose - - - - 26.1 12.4 71.4 34.3 RCSANet  [39] ICCV 21 RGB 48.6 50.2 100.0 97.2 - - - - 3DSL  [3] CVPR 21 RGB+pose+sil.+3D - 51.3 - - 31.2 14.8 - - FSAM  [2] CVPR 21 RGB+pose+sil. 54.5 - 98.8 - 38.5 16.2 73.2 35.4 GI-ReID  [5] CVPR 22 RGB+sil. - 37.5 - - 23.7 10.4 63.2 29.4 BSGA+CRE  [15] BMVC 22 RGB+parsing 61.8 58.7 99.6 97.3 - - - - CAL  [9] CVPR 22 RGB 55.2 55.8 100.0 99.8 40.1 18.0 74.2 40.8 CCFA  [41] CVPR 23 RGB 61.2 58.4 99.6 98.7 45.3 22.1 75.8 42.5 AIM  [42] CVPR 23 RGB 57.9 58.3 100.0 99.9 40.6 19.1 76.3 41.1 DCR-ReID  [40] TCSVT 23 RGB+parsing 57.2 57.4 100.0 99.7 41.1 20.4 76.1 42.3 chan et al.  [10] ACM 23 RGB 58.4 58.6 100.0 99.7 32.9 15.4 73.4 36.9 MADE RGB+description 64.3 59.1 100.0 98.6 47.4 24.4 84.2 48.2
@@
Results on PRCC. We compare our method with twelve cloth-changing re-identification methods on PRCC in Table I. We can notice that our method outperforms all other methods in the cloth-changing setting. Compared with AIM, the method of mining original RGB images in the cloth-changing setting, the rank-1 of our method increased by 6.4%, and the mAP increased by 0.8%. These shows that using multi-modal attribute description in CC-ReID to assist re-identification effectively improves results. Compared with the best method using multi-modal biometrics, BSGA+CRE, in the cloth-changing setting, the rank-1 increased by 2.5%, and the mAP increased by 0.4%. It shows that our method uses editable multi-modal attribute information and has better results when it is more convenient to remove cloth information than biological information. Data augmentation can significantly improve the model improvement effect, and our method is even better than CCFA, which uses feature enhancement. In the cloth-changing setting, the rank-1 increased by 3.1%, and the mAP increased by 0.7%.
@@
Results on LTCC. We compare our method with eight cloth-changing re-identification methods on LTCC in Table I. Our method outperforms all other methods in both cloth-changing and general settings. Compared with CCFA, in the cloth-changing setting, The rank-1 increased by 2.1%, and the mAP increased by 2.3%. In the general setting, The rank-1 increased by 8.4%, and the mAP increased by 5.7%. Compared with AIM, in the cloth-changing setting, The rank-1 increased by 6.8%, and the mAP increased by 5.3%. In the general setting, The rank-1 increased by 7.9%, and the mAP increased by 7.1%. Compared with chan et al., in the cloth-changing setting, The rank-1 increased by 14.5%, and the mAP increased by 9.0%. In the general setting, The rank-1 increased by 10.8%, and the mAP increased by 11.3%. It shows that our method embedding fused image features with attribute descriptions is better than using GAN networks to mine cloth-irrelated features of original images.
@@
TABLE II: Evaluations on Celeb-reID-light(%). Bold and underlined numbers are the top two scores.
@@
Method Type Method Venue Celeb-reID-light rank-1 mAP Traditional OSNet  [44] ICCV 19 21.3 11.7 DG-Net  [45] CVPR 19 23.5 12.6 BoT(resnet50)  [46] CVPRW 19 24.2 13.6 AGW(resnet50_nl)  [47] TPAMI 21 30.2 15.4 TransReID  [48] ICCV 21 31.3 18.6 CC-ReID RCSANet  [39] CVPR 21 29.5 16.7 MBUNet  [4] ICME 22 33.9 21.3 IRANet  [43] IVC 22 46.2 25.4 DeSKPro  [32] ICIP 22 52.0 29.8 MADE 72.0 52.3
TABLE III: Evaluations on LaST(%). Bold and underlined numbers are the top two scores.
@@
Method Type Method Venue LaST rank-1 mAP Traditional OSNet  [44] ICCV 19 64.3 21.0 BoT  [46] CVPRW 19 67.1 23.6 HOReID  [49] CVPR 20 68.3 25.5 Top-DB-Net  [50] ICPR 20 69.4 25.0 CtF  [51] ECCV 20 70.0 26.5 CC-ReID CAL  [9] CVPR 22 73.7 28.8 MADE 79.0 40.9
@@
Results on Celeb-reID-light and LaST. We compare our method with some traditional and cloth-changing re-identification methods on the two datasets in Table II and Table III. Our method outperforms all previous methods. The face is the most direct information for re-identification. For Celeb-reID-light, our method is even better than DeSKPro, which uses facial features. The rank-1 increased by 20.0%, and the mAP increased by 22.5%. For LaST, our method is better than CAL. The rank-1 increased by 5.3%, and the mAP increased by 12.1%. LaST is a large and challenging dataset that requires high model training complexity. Currently, there are few CC-ReID methods tested using LaST. It shows that our method is effective, has low model complexity, and can achieve testing of large datasets.
IV-E Ablation Study
@@
We use EVA02-large  [13] as the baseline and also use the loss function in section III-D for supervision. In this section, we explore the role of masked attribute description on the model’s learning of clothes-irrelevant features and the impact of different layer numbers in EVA02-large.
IV-E1 Effectiveness of Mask Attribute and Influence of Attribute Detection Accuracy
@@
We embedded the attribute description vector after masking cloth information into the baseline. Considering the accuracy of the attribute extraction model, we also introduced a certain proportion of random 0-1 noise into the masked attribute and summarized the experimental results in Table IV. To verify the effectiveness of attribute description in improving re-identification results. We conducted experiments on the cloth-changing setting of PRCC, LTCC and Celeb-reID-light. In all datasets, the performance of adding masked attribute descriptions is almost higher than the baseline.
The results of adding masked-attribute descriptions were higher than the baseline. After masking the cloth information in the attribute description without noise and embedding, for PRCC, the Rank-1 increased by 4.2%, and the mAP increased by 5.2%. For LTCC, the Rank-1 increased by 2.3%, and the mAP increased by 1.5%. Considering the accuracy issues of the attribute extraction model, we introduced random noise of 5%, 10%, 15%, and 20% separately into the attribute descriptions of every person in the PRCC, LTCC, and Celeb-reID-light datasets. Due to the varying effects of different noise levels on the improvement of attribute description data across different datasets, we report the average results of the experiments. When the noise level was 10%, the average rank-1 value across the three datasets was the highest, so we selected the experimental results with 10% noise as the final result. Specifically, compared to the baseline, when 10% noise was introduced, the rank-1 for PRCC increased by 2.3%, for LTCC increased by 3.8%, and for Celeb-reID-light increased by 4.2%. The model is more accurate in person re-identification, indicating that this method can compel the model to discard cloth information and learn cloth-insensitive features.
Then we discuss the impact of the accuracy of attribute detection models on experimental results. The attribute detection model we used is SOLIDER [12]. It has achieved excellent performance on widely used pedestrian attribute recognition datasets PETA_ZS [34], RAP_ZS [34], and PA100K [52], with mean accuracy (mA) of 76.4, 76.4, and 86.4, respectively [12]. Although its accuracy on the attribute recognition dataset did not achieve complete correctness, our method is robust to a certain proportion of errors in attribute detection. Since the cloth-changing dataset lacks attribute labels, it is not feasible to directly measure the accuracy of attribute recognition. Hence, we introduce random noise of 5%, 10%, 15%, and 20% separately into the attribute description of every person in PRCC, LTCC, and Celeb-reID-light. The aim is to investigate the influence on person re-identification results when the attribute recognition model is not sufficiently accurate. Following the introduction of noise, the experimental results of PRCC and LTCC under the cloth-changing setting and Celeb-reID-light are shown in Table IV.
In the LTCC dataset, appropriately adding random noise can improve re-identification accuracy, but adding more than 10% noise leads to a slight deterioration in results. However, in the PRCC dataset, adding a certain proportion of noise generally leads to a slight decrease in re-identification results, while adding 20% noise increases the rank-1 by 5.7% compared to the baseline. One possible reason is the difference in dataset styles. As shown in Fig. 6, pedestrian images in LTCC are generally darker overall, resulting in lower attribute recognition accuracy. Adding appropriate random noise can enhance pedestrian attribute recognition performance. On the other hand, images in PRCC have more apparent colors and pedestrian attributes are relatively easier to identify than those in LTCC. Therefore, adding a certain proportion of random noise decreases recognition performance. SOLIDER training uses PETA_ZS, a dataset collected from real-world scenarios, as shown in Fig. 6(a), whereas Celeb-reID-light is collected from the internet, as shown in Fig. 6(d). The difference in dataset styles may lead to higher re-identification performance when noise is added to Celeb-reID-light compared to when no noise is added.
In addition, Table V compares our model with the baseline on the PRCC dataset in terms of experimental scale. The experiment follows the setup described in Section IV-C. We calculated the time required for the model to train one epoch (batch size = 8), the FLOPs for a single image input to the model, and the model’s parameters. During the testing phase, we removed the attribute description information and calculated the time required for testing with a batch size of 128.
@@
TABLE IV: Ablation studies of attribute description of MADE in cloth-changing setting on PRCC, LTCC and Celeb-reID-light. Where m-att means masked attribute
Method PRCC LTCC Celeb-reID-light Average rank-1 mAP rank-1 mAP rank-1 mAP rank-1 mAP baseline 62.0 57.8 43.6 23.5 65.4 46.8 57.0 42.7 baseline w/ m-att (0% noise) 66.2 63.0 45.9 25.0 67.3 47.3 59.8 45.1 baseline w/ m-att (5% noise) 65.7 61.5 46.4 23.3 70.9 51.1 61.0 45.3   baseline w/ m-att (10% noise) 64.3 59.1 47.4 24.4 72.0 52.3 61.2 45.3 baseline w/ m-att (15% noise) 65.6 60.9 42.9 21.1 69.6 48.2 59.4 43.4 baseline w/ m-att (20% noise) 67.7 64.7 40.1 19.0 67.4 47.3 58.4 43.7
Figure 6: (a) Examples of PETA_ZS. (b) Examples of PRCC. (c) Examples of LTCC. (d) Examples of Celeb-reID-light. PETA_ZS, PRCC, and LTCC are datasets collected from real-world scenarios. Photos of LTCC exhibit an overall dark style, which may adversely affect the accuracy of attribute recognition models. PRCC dataset features a bright style that facilitates the identification of pedestrian attributes. Celeb-reID-light is a dataset collected from the internet.
TABLE V: The comparison of the experimental scale between the Baseline and our model on the PRCC dataset.
Method Training Testing Epoch time FLOPs Params CPU time FLOPs Params baseline 7m42s 162.4GFlops 304.1M 6h0m 162.4GFlops 304.1M ours 10m20s 162.8GFlops 312.8M 6h0m 162.8GFlops 312.8M
IV-E2 Gradually Masking cloth-Related Attributes
In this chapter, we gradually mask cloth-related attributes to validate our motivation. To prevent any specific attribute from influencing person re-identification results, we randomly mask these attributes at 30%, 60%, and 90%, progressing up to 100%, as depicted in Table VI. Experiments are conducted in the cloth-changing setting of PRCC and LTCC datasets. When 60% of the cloth-related attributes are masked, compared to 30%, the rank-1 of PRCC increases by 0.9%, and the rank-1 of LTCC increases by 1.3%. When 90% of the cloth-related attributes are masked, the mAP of PRCC increases by 2.9%, and the mAP of LTCC increases by 0.3%. When 100% of the cloth-related attributes are masked, the rank-1 of PRCC increases by 4.0%, and the rank-1 of LTCC increases by 2.3%. As the cloth-related attributes are gradually masked, the accuracy of re-identification improves. The re-identification accuracy remains relatively high even when the cloth-related attributes are partially masked. This suggests that attributes related to clothes affect model re-identification results, and their removal forces the model to learn cloth-insensitive features.
TABLE VI: The experimental results of progressively masking pedestrian cloth-related attributes in the cloth-changing setting of MADE on PRCC and LTCC.
masking ratio PRCC LTCC rank-1 mAP rank-1 mAP 30% 62.2 59.4 43.6 22.1 60% 63.1 59.4 44.9 21.4 90% 66.1 62.3 44.9 22.4 100% 66.2 63.0 45.9 25.0
IV-E3 Number of Stages and TrV Blocks